# quantum neural networks, Bernstein’s inequality, neural tangent kernel, Clifford group

### v1

@@ 5 Conclusions
We have proposed an efficient classical algorithm to estimate the NTK arising from a broad family of parametric quantum circuits constructed using Clifford unitaries that depend arbitrarily on the input interleaved by parametric Pauli rotations. Despite the parametric Pauli rotations belong to the Clifford group only for few special values of the parameters, we demonstrated that the proposed Algorithm 1 can always efficiently estimate the NTK of the model with provable accuracy. Moreover, we have showed that Algorithm 1 can be combined with the recently proved equivalence between very wide trained quantum neural networks and Gaussian processes girardi2024 ; hernandez2024 to efficiently estimate the mean of the trained model function in the limit of infinite width (Algorithm 2). Therefore, our results prove that quantum neural networks cannot have quantum advantage in such a limit.
Our work aligns with a recent series of results ReardonSmith2024improvedsimulation ; PRXQuantum.6.010337 ; martinez25 ; cerezo2024doesprovableabsencebarren ; angrisani2024 showing that the output of several parametric quantum circuits can be estimated efficiently using classical resources. These findings suggest a need for a more refined understanding of where true quantum advantage lies, and under what architectural constraints an efficient classical simulation is feasible.
@@ Appendix A Bernstein’s upper bound for random matrices
Theorem 6 (Matrix Bernstein (tropp2015, , Theorem 6.1.1)).

### v2

@@ 5 Conclusions
We have proposed an efficient classical algorithm to estimate the NTK arising from a broad family of parametric quantum circuits constructed using Clifford unitaries that depend arbitrarily on the input interleaved by parametric Pauli rotations. Despite the parametric Pauli rotations belong to the Clifford group only for few special values of the parameters, we demonstrated that the proposed Algorithm 1 can always efficiently estimate the NTK of the model with provable accuracy. Moreover, we have showed that Algorithm 1 can be combined with the recently proved equivalence between very wide trained quantum neural networks and Gaussian processes [11, 20] to efficiently estimate the mean of the trained model function in the limit of infinite width (Algorithm 2). Therefore, our results prove that wide quantum neural networks with logarithmic depth trained on supervised-learning problems with the input encoded with Clifford gates can be simulated efficiently by classical computers and cannot achieve quantum advantage.
Our work aligns with a recent series of results [22, 33, 18, 7, 3] showing that the output of several parametric quantum circuits can be estimated efficiently using classical resources. These findings suggest a need for a more refined understanding of where true quantum advantage lies, and under what architectural constraints an efficient classical simulation is feasible.
@@ Appendix A Bernstein’s upper bound for random matrices
Theorem 6 (Matrix Bernstein [31, Theorem 6.1.1]).